OpenStudy (anonymous):

USING THE ELIMINATION METHOD,SOLVE THE SYSTEM OF EQUATIONS: A+9B=11 A+2B= -3

OpenStudy (unklerhaukus):

take the second equation from the first

OpenStudy (unklerhaukus):

solve for B, then back substitute to find A

OpenStudy (unklerhaukus):

\[A+9B=11\qquad\qquad{\small(i)}\]\[A+2B= -3\qquad\qquad{\small(ii)}\] \[(A+9B)-(A+2B)=11-(-3)\qquad\qquad{\small(i)-(ii)}\]

OpenStudy (anonymous):

SO THE ANSWER IS 2,11

OpenStudy (unklerhaukus):

to check your answer plug into one of the original equations to see if B=2, A=11 is right ; \[A+9B=11\qquad\Longrightarrow 11+9\times2=11?\]

OpenStudy (anonymous):

B=2 &A =11 RIGHT

OpenStudy (unklerhaukus):

that dosent work

OpenStudy (anonymous):

RIGHT

OpenStudy (unklerhaukus):

\[(A+9B)-(A+2B)=11-(-3)\qquad\qquad{\small(i)-(ii)}\]\[A-A+9B-2B=11+3\]\[7B=14\]\[B=2\]

OpenStudy (unklerhaukus):

\[A+9\times(2)=11\]\[A=\dots?\]

OpenStudy (anonymous):

2

OpenStudy (unklerhaukus):

B is two , you have got this right, but what is A?

OpenStudy (unklerhaukus):

A=2,B=2, dosent work because \[2+9\times2=20\neq11\]

OpenStudy (anonymous):

SO IM LOST..

OpenStudy (unklerhaukus):

you have found B right?

OpenStudy (anonymous):

RIGHT

OpenStudy (anonymous):

but i thought the way i was doinit was correct...guess not :(

OpenStudy (unklerhaukus):

there are heaps of ways to solve this sort of problem, if you'd rather a different method thats fine but you'll have to let me know what it is, if you want some help we have \(B=2\) so looking back at the original equations \[A+9B=11\qquad \Rightarrow A+9(2)=11\] \[A+2B= -3\qquad \Rightarrow A+2(2)= -3\] all we need is A

OpenStudy (unklerhaukus):

Using the first equation we could find A like this:\[A+9\times2=11\]\[A+18=11\]\[A=11-18\] Or using the second equation \[A+2\times2=−3\]\[A+4=-3\]\[A=-3-4\]

OpenStudy (unklerhaukus):

which ever method you use you'll find that A is .....

OpenStudy (anonymous):

-7

OpenStudy (unklerhaukus):

if you are right both these should be true (ie LHS=RHS) (-7)+9(2)=11 (-7)+2(2)= -3

OpenStudy (unklerhaukus):

what are the left hand sides (LHS's)

OpenStudy (unklerhaukus):

-7+9x2=-7+18=... -7+2x2=-7+4=... /?

OpenStudy (unklerhaukus):

what do you get? dose A=-7, B=2 work for both of the original equations?

OpenStudy (unklerhaukus):

if so, you have solved the system